Suppose that your company uses a continuous review system, and you have an average monthly demand of 12,000 units with a monthly standard deviation of 400 units. The time it takes for your supplier to deliver to your plant is 5 days (assume 30 days per month). If your company has a CSL level of 50%.

  1. How much safety stock should be held?
  2. Now suppose that the company is considering switching to a periodic review system, and they plan to review their inventory position every 7 days. Assuming the rest of the information given above remains the same, how much safety stock should be held?
  3. If the company wanted to maintain the same level of safety stock in the periodic review system as they had in the continuous review system at what cycle service level would they be operating?
Answer #1

Monthly demand, d = 12000 units

Standard deviation, s = 400 units

Lead time, L = 5 days = 5/30 months

CSL = 50%

z = NORMSINV(.50) = 0

Safety stock required = z*s*sqrt(L)

= 0*400*sqrt(5/30)

= 0 units

b) Review period, P = 7 days

Protection interval = P+L = 5+7 = 12 days = 12/30 = 0.4 month

Safety stock required = z*s*sqrt(P+L)

= 0(400*sqrt(0.4)

= 0 units

c) For 50% CSL, continuous review system and periodic system both have the same safety stock level.